# Compelling ReLU Network Initialization and Training to Leverage Exponential Scaling with Depth

### v3

@@ 3.1 Differentiable Model Output
Given the rapid oscillations of the triangle waves formed at each layer, the resulting output of the network will appear jagged with most choices of scaling parameters. This can lead to the network settling into overfit local minima, in which the data points might lie along the curve, but the curve is not predictive of the areas in between due to rapid oscillation. Our main mathematical result addresses this issue by forcing differentiability of the network output. It turns out that for this class of functions, the peak locations of the triangle waves uniquely determine the scales with which to sum them.
@@
4 Experiments
The goal of deep learning is to train a network to approximate a (generally unknown) nonlinear function. Accordingly, in this section, we implement our method to learn several convex one-dimensional curves. These nonlinear functions are known—so that we can measure error—and we remark on more difficult functions in Sections 5 and 6. The aim of these experiments is twofold: (1) we would like to determine how to learn the most effective function representations possible, and (2) to explore how the utilization of an increased number of linear regions can affect a network’s ability to capture underlying nonlinearity in its training data. To demonstrate that our networks can learn function representations that better utilize depth, we benchmark against PyTorch’s default settings (nn.linear() uses Kaiming initialization), as well as the RAAI distribution from Singh and Sreejith (2021), and produce errors that are orders of magnitude lower than both. For our experimental networks, we use a common set of initialization points where the triangle peaks and scaling parameters are chosen according to our main theorem (Equation (4)). We compare the effect of reparameterized pretraining against skipping immediately to standard gradient descent. We also compare training the scaling parameters freely during pretraining, instead of choosing them to achieve differentiability. The intent of both comparisons is see if pretraining and differentiability constraints facilitate smoother navigation of the loss landscape and avoidance of local minima. Lastly, we conduct a second round of tests to determine if pretrained networks display an enhanced predictive capacity on unseen data points, as might be expected if they can leverage greater nonlinearity in their outputs.
@@ 4.1 Experimental Setup
All models are trained using Adam (Kingma and Ba, 2017) as the optimizer with a learning rate of 0.0001 for 1,000 epochs to ensure convergence. Each network is four neurons wide with five hidden layers, along with a one-dimensional input and output. The loss function used is the mean squared error, and the average and minimum loss are recorded for 30 models of each type. The networks unique to this paper share a common set of starting locations, so that the effects of each training regimen are directly comparable. As in related papers (Perekrestenko et al., 2018; Daubechies et al., 2022), we restrict our attention to one-dimensional examples, which are sufficient to demonstrate our proposed theory and methodology. The four curves we approximate are $x^{3}$, $x^{11}$, $\tanh(3x)$, and a quarter period of a sine wave. The curves are chosen to capture a variety of convex one-dimensional functions. To approximate the sine and the hyperbolic tangent, the triangle waves are added to the line $y=x$. For the other approximations, the waves are subtracted. This requires the first scaling factor to be $a_{0}*a_{1}$ instead of $(1-a_{0})*a_{1}$. The first set of data is 500 evenly spaced points on the interval $[0,1]$ for each of the curves. This is chosen to be very dense deliberately, to try to evoke the most accurate representations from each network. We determine from these tests that pretraining with differentiability enforced produces the best results, so we compare it to standard networks in our second set of experiments. We use a second set of data consisting of only 10 points, with a test set of 10 points spaced in between so as to be as far away from learned data as possible. The goal of this set of experiments is to compare the predictive capacities of the networks on unseen data.
@@ 4.2 Numerical Results
Pretraining with differentiability enforced also closes the gap between the minimum and mean errors compared to other setups. This indicates that these loss landscapes are indeed the most reliable to traverse. Even in our noiseless setting, we find that it is still possible to settle into local minima where the function fits only a small subset of the data points and fails to accurately represent the target curve in between them. Enforcing differentiability during pretraining can impart a bias towards smoother solutions during gradient descent and eliminates these occurrences in our experiments.
@@ 4.3 Gradient Descent does not Directly Optimize Efficiency
Figure 5: Approximation produced with standard methodologies. The layers are shown before ReLU is applied. The neuron colors here are arbitrary; they do not correspond to Figure 2.
Figure 5 shows the interior of a default network. The layers here are shown before applying ReLU. The default networks fail to make efficient use of ReLU to produce linear regions, even falling short of 1 bend per neuron, which can easily be attained by forming a linear spline (1 hidden layer) that interpolates some of the data points. Examining the figure, the first two layers are wasted. No neuron’s activation pattern crosses $y=0$, so ReLU is never used. Layer 3 could be formed directly from the input signal. Deeper in the network, more neurons remain either strictly positive or negative. Those that intersect $y=0$ are monotonic, only able to introduce one bend at a time. The core issue is that while more bends leads to better accuracy, networks that have few bends are not locally connected in parameter space to those that have many. This is problematic since gradients can only carry information about the effects of infinitesimal parameter modifications. If a bend exists, gradient descent can reposition it. But for a neuron that always outputs a strictly positive value (such as the red in layer 2), bends cannot be introduced by infinitesimal weight or bias adjustments. Therefore, bend-related information will be absent from its gradients. Gradient descent will only compel a network to bend by happenstance; indirectly related local factors must guide a neuron to begin outputting negative values. Rarely, these local incentives are totally absent, and the network outputs a bend-free line of best fit.
4.4 Effect of Pretraining on a Compositional Network
@@
In Figure 6 we compare the top performing models with and without reparameterized pretraining. We observe that without the guidance of the pretraining, gradient descent usually loses the triangle generating structure. It typically happens around layer 4 or 5, devolving into noisy patterns and resulting in higher errors, whereas pretraining maintains structure at greater depths. This behavior of gradient descent we observe is problematic since theoretical works often rely on specific constructions within networks to prove their results. Gradient descent greedily abandons any such structure in favor of models that can be worse in the long term. A theoretical result that shows a certain representation exists in the set of neural networks will thus have a hard time actually learning it without a subsequent plan to control training.
5 Limitations and Extensions
Although our reparemeterization can help maintain the triangular structure out to greater depths than unassisted gradient descent, our neurons will still begin producing non-triangular outputs eventually. This suggests that our reparameterization is limited in the functions it can theoretically capture. It may be enabling a rapid error reduction initially in the first few layers, but be unable to approach most curves to arbitrary precision. It is in this remaining error gap that the second stage, gradient descent, becomes lost. In the appendix, we detail two important results concerning the theoretical expressive power of our reparameterization that support this.
First, the differentiability constraints on the scaling parameters during pretraining necessitate that the network outputs a convex function. It is important for the pretraining to be able to produce a reasonably close initial solution; therefore, the reparameterization we use is only guaranteed to be suitable for approximating convex functions. The second limitation we encounter is that pursuing further smoothness (at least in this 4-neuron configuration) is not possible. The appendix show that approximating $x^{2}$ is the only way to have a twice continuously differentiable output in the infinite depth limit.
Despite these limitations, these networks could still be considered as replacements of the $x^{2}$ gate discussed in Section 2.1. Likewise, they could be used as a base component for assembly into larger networks. These composed networks could learn non-convex functions in higher dimensions, and they would at minimum be guaranteed to represent interpolating polynomials. The drawback is that they would likely become prohibitively large. Rather than approximating another approximation scheme out of ReLU networks, it may be better to focus on finding representations that more naturally arise from ReLU networks themselves when extending this work. Two properties specifically stand out as being somewhat fundamental to deep ReLU approximations: each linear region produced is a full copy of the original input interval $[0,1]$, and the linear regions’ boundaries become dense as depth tends towards infinity. Density ensures that no two inputs are linearly related in an infinite network, and the repeated reproduction of the input space allows each layer to act on each previously existing region independently and simultaneously. These properties both arise in our method, and are likely important considerations when seeking theoretical extensions. Lastly, it might be important to consider non-differentiable model outputs as a way to gain additional flexibility, but additional protections would need to be put in place against overfitting.

### v4

@@ 3.1 Differentiable Model Output
Given the rapid oscillations of the triangle waves formed at each layer, the resulting output of the network will appear jagged or fractal with many choices of scaling parameters. This can lead to the network settling into solutions that are not predictive of the areas in between training points (or possibly not even interpolating the training data) due to rapid oscillation. Our main mathematical result addresses this issue by forcing differentiability of the network output. To achieve differentiability, it turns out that for this class of functions, the peak locations of the triangle waves uniquely determine the scales with which to sum them:
@@
4 One-Dimensional Experiments
The goal of deep learning is to train a network to approximate a (generally unknown) nonlinear function. Accordingly, in this section, we implement our method to learn several convex one-dimensional curves. These nonlinear functions are known—so that we can measure error—and we remark on more difficult functions in Sections 5 and 6. The aim of these experiments is twofold: (1) we would like to determine how to learn the most effective function representations possible, and (2) to explore how the utilization of an increased number of linear regions can affect a network’s ability to capture underlying nonlinearity in its training data. To demonstrate that our networks can learn function representations that better utilize depth, we benchmark against PyTorch’s (Paszke et al., 2019) default settings (nn.linear() uses Kaiming initialization), as well as the RAAI distribution from Singh & Sreejith (2021), and produce errors that are orders of magnitude lower than both. For our experimental networks, we use a common set of initialization points where the triangle peaks and scaling parameters are chosen according to our main theorem (Equation 4). We compare the effect of reparameterized pretraining against skipping immediately to standard gradient descent (from a reparameterized initialization as in Section 3). We also compare training the scaling parameters freely during pretraining, instead of choosing them to achieve differentiability. The intent of both comparisons is see if pretraining and further differentiability constraints facilitate smoother navigation of the loss landscape. Lastly, we conduct a second round of tests to determine if pretrained networks display an enhanced predictive capacity on unseen data points, as might be expected if they can leverage greater nonlinearity in their outputs.
@@ 4.1 Experimental Setup
All models are trained using Adam (Kingma & Ba, 2017) as the optimizer with a learning rate of 0.001 for 1,000 epochs to ensure convergence (see Appendix for a brief analysis of different learning rates). Each network is four neurons wide with five hidden layers, along with a one-dimensional input and output. The loss function used is the mean squared error, and the average and minimum loss are recorded for 30 models of each type. The networks unique to this paper share a common set of starting locations using the construction in Section 3, so that the effects of each training regimen are directly comparable. As in related papers (Perekrestenko et al., 2018; Daubechies et al., 2022), we focus for the moment on one-dimensional examples, which are sufficient to demonstrate our proposed theory and methodology. The four curves we approximate are $x^{3}$, $x^{11}$, $\tanh(3x)$, and a quarter period of a sine wave. The curves are chosen to capture a variety of convex one-dimensional functions. To approximate the sine and the hyperbolic tangent, the triangle waves are added to the line $y=x$. For the other approximations, the waves are subtracted. This requires the first scaling factor to be $a_{0}*a_{1}$ instead of $(1-a_{0})*a_{1}$. The first set of data is 500 evenly spaced points on the interval $[0,1]$ for each of the curves. This is chosen to be very dense deliberately, to try to evoke the most accurate representations from each network. We determine from these tests that pretraining with differentiability enforced produces the best results, so we compare it to standard networks in our second set of experiments. We use a second set of data consisting of only 10 points, with a test set of 10 points spaced in between so as to be as far away from learned data as possible. The goal of this set of experiments is to compare the predictive capacities of the networks on unseen data.
@@ 4.2 Numerical Results
Pretraining with differentiability enforced also closes the gap between the minimum and mean errors compared to other setups. This indicates that these loss landscapes are indeed the most reliable to traverse. Enforcing differentiability during pretraining can impart a bias towards smoother solutions during subsequent unassisted gradient descent.
@@ 4.3 Gradient Descent does not Directly Optimize Efficiency
Figure 5: Approximation produced with standard methodologies (shown before ReLU is applied). The neuron colors here are arbitrary; they do not correspond to Figure 2.
Figure 5 shows the interior of a default network. The layers here are shown before applying ReLU. The default networks fail to make efficient use of ReLU to produce linear regions, even falling short of 1 bend per neuron, which can easily be attained by forming a linear spline (1 hidden layer) that interpolates some of the data points. Examining the figure, the first two layers are wasted. No neuron’s activation pattern crosses $y=0$, so ReLU is never used. Layer 3 could be formed directly from the input signal. Deeper in the network, more neurons remain either strictly positive or negative. Those that intersect $y=0$ are monotonic, only able to introduce one bend at a time. The core issue is that while more bends leads to better accuracy, networks that have few bends are not locally connected in parameter space to those that have many. This is problematic since gradients can only carry information about the effects of infinitesimal parameter modifications. If a bend exists, gradient descent can reposition it. But for a neuron that always outputs a strictly positive value (such as the red in layer 2), bends cannot be introduced by infinitesimal weight or bias adjustments. Therefore, bend-related information will be absent from its gradients. Gradient descent will only compel a network to bend by happenstance; indirectly related local factors must guide a neuron to begin outputting negative values. Occasionally, these local incentives are totally absent, and the network outputs a bend-free line of best fit.
4.4 Effect of Reparameterized Initialization vs Pretraining
@@
Figure 6 compares the top performing reparameterized models for $x^{3}$ with and without pretraining. We observe that without the guidance of the pretraining, gradient descent usually loses the triangle generating structure around layer 4 or 5, devolving into noisy patterns and resulting in higher errors. Pretraining maintains structure at greater depths. This behavior of gradient descent we observe is problematic since theoretical works often rely on specific constructions within networks to prove their results. Gradient descent greedily abandons any such structure during its optimization in favor of models that can be worse in the long term. A theoretical result that shows a certain representation exists in the set of neural networks will thus have a hard time actually learning it without a subsequent plan to control training. This phenomenon highlights the need for theoretical work to derive more expressive training controls, which can give better guidance for the optimization of functions of greater dimension or complexity.
5 Extension to Non-Convex Functions and Higher Dimensions
Figure 7: Approximation of $y=x^{3}-x$ by difference of pretrained components, achieving a loss of $4\times 10^{-7}$. A standard $8\times 5$ network yielded a larger loss of $8\times 10^{-6}$.
The results presented so far are limited to one-dimensional convex functions, but these are not limitations of our method. Our constructed networks can be applied to approximate higher-dimensional functions by using these networks as activation functions in a larger network (see Figure 9 in the Appendix for a schematic). For instance, in two dimensions, we take two or more copies of our network and estimate a target function as a linear combination of the networks’ outputs. Because the weights of the linear combination are randomized, each copy of the network will tend to learn different lower-dimensional aspects of the target function, whereby the combination can accurately capture the higher-dimensional function.
The same approach can address non-convex functions since many non-convex functions (including every continuous function with a bounded second derivative (Hessian)) can be expressed as the difference of two convex functions (see Zhang et al. (2018), which also describes how every ReLU network can be decomposed into a difference of two piecewise linear convex functions). We do not claim that two copies of our network are sufficient for learning every pathological function, but preliminary results in this section are surprisingly promising (and of course, assembling more copies of our network into a deeper or wider larger network would give additional representational power).
We note that the approach in this section bears similarity to the recently popularized Kolmogorov-Arnold Networks (KAN) (Liu et al., 2024), which use spline functions as activations of a network where every weight is 1. This is based on the Kolomogorov-Arnold representation theorem (Kolmogorov, 1957; Arnold, 1957; 1959), which gives a guarantee that every continuous multivariate function can be represented by a neural network-like structure with one-dimensional activation functions. But by restricting the activation functions to splines, the KAN abandons its theoretical guarantees. Using our networks as activations, however, maintains the theoretical ability to perform polynomial interpolation since they retain the ability to approximate $x^{2}$, and the multiplication gate discussed in Section 2.1 is based on a linear combination of squared terms.
Figures 7 and 8 show learning the non-convex function $y=x^{3}-x$ and the two-dimensional function $z=r^{3}$ on $[-1,1]$, each by using two of our convex function blocks. Rather than directly using our networks as activation functions, we instead encode each of their layers as a block diagonal structure in one combined matrix (which has the same effect). Two slight modifications are made to each of the component networks to better interact with ReLU. First, two neurons are used for the sum so that a negative component may be maintained in the case of $x^{3}-x$, which would otherwise be zeroed out by ReLU. Second, the triangular functions are generated upside-down, preventing ReLU from zeroing the function outside of the approximation region.
For both $x^{3}-x$ and $r^{3}$, we substantially outperform standard initialization and training procedures. Especially striking is the increased number of linear regions in Figure 8. Visible in Figure 7 is the gradient descent-induced destructuring discussed in Section 4.4. This likely occurs due to the block diagonal structure of the larger weight matrices, which results in many zeros for the second pass of gradient descent to fill in unassisted. This could be addressed either by limiting training only to the block diagonal entries, or by deriving a more advanced pretraining method that makes use of the full matrix. Note that since $x^{3}-x$ is twice differentiable we can write it as $(x^{3}-x+3x^{2})-3x^{2}$, a difference of functions convex on $[-1,1]$ obtained by finding a parabola based on trying to make the derivative of $x^{3}-x$ monotone. Note that this is not a unique decomposition.
Figure 8: Approximations of $z=\sqrt{x^{2}+y^{2}}^{3}$ using an $8\times 5$ regular network (left) and a union of two of our pretrained components (right). Losses are $1.5\times 10^{-4}$ and $3.5\times 10^{-6}$, respectively, demonstrating a nearly two orders of magnitude improvement using our techniques.